# Sensory Glove-Based Surgical Robot User Interface

### v2

@@ I INTRODUCTION
One potential area for further improvement and technological innovation in robotic surgery is the integration of modern sensory gloves into robotic platforms, which allow surgeons to intuitively control robotic arms with their hand movements without the need for additional manipulators that are customary for existing consoles. Furthermore, most end-effectors of surgical robot instruments, such as the Fenestrated Bipolar Forceps (FBF) and the Large Needle Driver (LND), are limited to the use of opposing blades due to the design of the console manipulators (as the last joint is controlled only by pinching two fingers); glove-based control could offer the surgeon additional dexterity while using these instruments.
@@ II-A System Architecture
The system integrates several components. The surgical robot platform comprises the first-generation da Vinci robot paired with the dVRK. One of the Patient Side Manipulators (PSMs) is equipped with the Fenestrated Bipolar Forceps (FBF) instrument and is subject to glove control. The arm is configured as in Fig. 1. The HTC Vive Tracker attached to the Manus Meta Prime 3 XR glove provides the pose of the hand, while the fingers are tracked by the glove. Instead of using image-based hand-tracking devices as in [9], we decided to use sensor gloves because the hand-tracking performance decreases in the presence of visual occlusion or changes in room lighting. These occur frequently in the OR, such as when other surgical team members use secondary screens to look at details in the scene. Moreover, sensory gloves offer vibrotactile feedback. In addition, classifier models were trained for hand gesture recognition; hand gestures were implemented to substitute the da Vinci console’s pedal inputs. SCOPEYE smart glasses that provide the same visual feedback as the High-Resolution Stereo Video (HRSV) of the da Vinci console and have a maximum latency of 7ms [16] were used to further simplify the setup.
@@
For the full da Vinci system, the kinematic chain from the Patient Side Manipulator (PSM) tip to the Endoscope Camera Manipulator (ECM) tip involves Setup Joints (SUJs). The poses provided by the dVRK exhibit errors within a range of ±$\pm 5cm$ for translation and fall between $5\sim 10$ degrees for orientation. These inaccuracies arise due to challenges in calibrating the SUJs precisely, as reported in [15]. Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [20] that eliminates the influence of SUJs and allows us to accurately determine the transformation between the PSM base (frame $R$) and the instrument tip (frame $T$), see Fig. 1.
@@ II-F Clutch
TABLE II: Examples of point clouds used for gesture prediction.
@@ III-B Training Results
Model Accuracy F1 Score Recall
FNN 0.9857 0.9764 0.9784
L-GBM 0.9875 0.9873 0.9875
RF 0.9830 0.9826 0.9830
HGBT 0.9792 0.9789 0.9792
TABLE III: Evaluation metrics of the trained models.
Table III shows the evaluation results on the test set. Based on the metrics, all models had analogous performance in estimating the hand gestures. Hence, we conducted an additional live test in which the models predicted the gestures in real time for 1000 iterations each. As shown in Table IV, L-GBM was the fastest among the models with an average delay of 1.63 ms. The speed is mainly due to its efficient way of growing trees and making their structure as compact as possible. The hyperparameters that were used for training the current L-GBM model are as follows: learning_rate=0.1, max_depth=15, min_child_samples=10, n_estimators=100, and num_leaves=15.
@@
Model Mean Time (ms) Std. Deviation (ms)
FNN 2.70 0.95
L-GBM 1.63 1.59
RF 3.90 0.62
HGBT 4.54 1.40
TABLE IV: Real-time prediction time for each model (1000 trials).
@@ IV-B Control Results
TABLE V: The average translational and rotational errors between each controller and the PSM across all participants. The last column shows the average delay between the controller and the PSM tip signals.
@@
TABLE VI: Performance evaluation. The first column displays the average runtime, and the subsequent column shows the average traveled Euclidean distance of the instrument tip manipulated by each controller of the five trials from each participant.
To compare the input motion of the glove and the output motion of the instrument tip, we first reversed the scaling (Eq. (1)) of the movement of the instrument tip (Fig. 3). The figure shows a delay between the two trajectories, mainly due to the computation time for the inverse kinematics. To accurately estimate the error between the input and output signals, we identified the delay by the peak positions within each signal and eliminated the effect of the delay by shifting the PSM signal to align with the glove’s signal (Fig. 3). The difference in translation was quantified using the Euclidean distance, while the discrepancy in rotation was determined using the standard bi-invariant metric on the group of rotations [34]. The errors for each controller in all the trials across the participants are listed in Table V. Although the magnitude of the errors is larger for our system, the task completion times in the two setups were comparable (Table VI), suggesting that the errors were acceptable. These larger errors might also explain the slightly higher average travel distances for our system.
@@
We evaluated the average performance of the participants with each controller (Table VI). Performance is evaluated by the average duration and travel distance of the instrument tip, indicating how efficiently they managed to perform the task. Considering that this is a prototype system, the performance of controlling the robot with the glove compared favorably with that of the console.
@@ IV-C Survey Results
Participants were asked to compare the usability and task load of our proposed system with the conventional console. Usability was assessed with the System Usability Scale (SUS) [17], and task load was measured using NASA Task Load Index (TLX) [18] questionnaires. The overall results are shown in Table VII.
The average SUS score for the glove was 67.5 compared to the console’s 81.4 (13.9 points of difference). These results align well with the quantitative assessment of the performance in Table VI: users who achieved similar performance between the glove and the console responded with a similar SUS assessment for the two. Considering that users have significantly more training on the dVRK console and that our system is just a prototype, we judge the reported SUS score as overwhelmingly positive.
@@ V Conclusion
TABLE VII: System Usability Scale (SUS) and NASA Task Load Index (TLX) questionnaire results, answered by surgeons (green letters) and medical doctors (red letters). In the SUS, a score closer to 100 indicates better usability, and in the TLX, a score closer to 100 indicates a higher workload.
@@
The existing surgical consoles are bulky and take up valuable space in the operating room. They limit the ability to optimize the operating room setup for throughput. They also pose challenges to the communication of the surgical team, leading to potential workflow disruptions. In contrast, our proposed system is inexpensive, as it uses off-the-shelf components, and it dramatically reduces space requirements. It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field. Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
Future work includes implementing velocity control to enhance the user experience, enabling full control of both surgical arms,upgrading hardware to improve tracking precision, and expanding the repertoire of user gestures to increase the range of available actions.

### v3

@@ I INTRODUCTION
One potential area for further improvement and technological innovation in robotic surgery is the integration of modern sensory gloves into robotic platforms, which allow surgeons to intuitively control robotic arms with their hand movements without the need for additional manipulators that are customary for existing consoles. Furthermore, most end-effectors of surgical robot instruments, such as the Fenestrated Bipolar Forceps (FBF) and the Large Needle Driver (LND), are limited to the use of opposing blades  [10] due to the design of the console manipulators (as the last joint is controlled only by pinching two fingers); glove-based control could offer the surgeon additional dexterity while using these instruments.
@@ II-A System Architecture
The system integrates several components. The surgical robot platform comprises the first-generation da Vinci robot paired with the dVRK. One of the Patient Side Manipulators (PSMs) is equipped with the Fenestrated Bipolar Forceps (FBF) instrument and is subject to glove control. The arm is configured as in Fig. 1. The HTC Vive Tracker attached to the Manus Meta Prime 3 XR glove provides the pose of the hand, while the fingers are tracked by the glove. Instead of using image-based hand-tracking devices as in [9]. We decided to use sensory gloves because video-based hand-tracking (commonly used in AR, VR, and Leap Motion applications) performs poorly in the presence of visual occlusion, background noise, or variations in room lighting [20]. These occur frequently in the OR, such as when other surgical team members use secondary screens to look at details in the scene. Moreover, sensory gloves offer vibrotactile feedback. In addition, classifier models were trained for hand gesture recognition; hand gestures were implemented to substitute the da Vinci console’s pedal inputs. SCOPEYE smart glasses that provide the same visual feedback as the High-Resolution Stereo Video (HRSV) of the da Vinci console and have a maximum latency of 7ms [17] were used to further simplify the setup.
@@
For the full da Vinci system, the kinematic chain from the Patient Side Manipulator (PSM) tip to the Endoscope Camera Manipulator (ECM) tip involves Setup Joints (SUJs). The poses provided by the dVRK exhibit errors within a range of ±$\pm 5cm$ for translation and fall between $5\sim 10$ degrees for orientation. These inaccuracies arise due to challenges in calibrating the SUJs precisely, as reported in [16]. Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [22] that eliminates the influence of SUJs and allows us to accurately determine the transformation $g_{rt}$ between the PSM base (frame $R$) and the instrument tip (frame $T$), see Fig. 1.
@@ II-F Clutch
TABLE II: Examples of point clouds used as input
@@ III-B Training Results
Model Accuracy F1 Score Recall Mean (ms) Std. Dev. (ms)
FNN 0.9857 0.9764 0.9784 2.70 0.95
L-GBM 0.9875 0.9873 0.9875 1.63 1.59
RF 0.9830 0.9826 0.9830 3.90 0.62
HGBT 0.9792 0.9789 0.9792 4.54 1.40
TABLE III: Evaluation metrics of the trained models. The last two columns represent the real-time prediction time for each model (estimated over 1000 trials).
Table III shows the evaluation results on the test set. Based on the metrics, all models had analogous performance in estimating the hand gestures. Hence, we conducted an additional live test in which the models predicted the gestures in real time for 1000 iterations each. As shown in Table III, L-GBM was the fastest among the models with an average delay of 1.63 ms. The speed is mainly due to its efficient way of growing trees and making their structure as compact as possible. The hyperparameters that were used for training the current L-GBM model are as follows: learning_rate=0.1, max_depth=15, min_child_samples=10, n_estimators=100, and num_leaves=15.
@@ IV-B Control Results
TABLE IV: The average translational and rotational errors between each controller and the PSM across all participants. The last column shows the average delay between the controller and the PSM tip signals.
@@
TABLE V: Performance evaluation. The first column displays the average runtime, and the subsequent column shows the average traveled Euclidean distance of the instrument tip manipulated by each controller of the five trials from each participant.
To compare the input motion of the glove and the output motion of the instrument tip, we first reversed the scaling (Eq. (1)) of the movement of the instrument tip (Fig. 3). The figure shows a delay between the two trajectories, mainly due to the computation time for the inverse kinematics. To accurately estimate the error between the input and output signals, we identified the delay by the peak positions within each signal and eliminated the effect of the delay by shifting the PSM signal to align with the glove’s signal (Fig. 3). The difference in translation was quantified using the Euclidean distance, while the discrepancy in rotation was determined using the standard bi-invariant metric on the group of rotations [36]. The errors for each controller in all the trials across the participants are listed in Table IV. Although the magnitude of the errors is larger for our system, the task completion times in the two setups were comparable (Table V), suggesting that the errors were acceptable. These larger errors might also explain the slightly higher average travel distances for our system.
@@
We evaluated the average performance of the participants with each controller (Table V). Performance is evaluated by the average duration and travel distance of the instrument tip, indicating how efficiently they managed to perform the task. Considering that this is a prototype system, the performance of controlling the robot with the glove compared favorably with that of the console.
@@ IV-C Survey Results
Participants were asked to compare the usability and task load of our proposed system with the conventional console. Usability was assessed with the System Usability Scale (SUS) [18], and task load was measured using NASA Task Load Index (TLX) [19] questionnaires. The overall results are shown in Table VI.
The average SUS score for the glove was 67.5 compared to the console’s 81.4 (13.9 points of difference). These results align well with the quantitative assessment of the performance in Table V: users who achieved similar performance between the glove and the console responded with a similar SUS assessment for the two. Considering that users have significantly more training on the dVRK console and that our system is just a prototype, we judge the reported SUS score as overwhelmingly positive.
@@ V Conclusion
TABLE VI: System Usability Scale (SUS) and NASA Task Load Index (TLX) questionnaire results, answered by surgeons (green letters) and medical doctors (red letters). In the SUS, a score closer to 100 indicates better usability, and in the TLX, a score closer to 100 indicates a higher workload.
@@
The existing surgical consoles are bulky and take up valuable space in the operating room. They limit the ability to optimize the operating room setup for throughput. They also pose challenges to the communication of the surgical team, leading to potential workflow disruptions. In contrast, our proposed system is inexpensive (estimated total cost of  \$5000), as it uses off-the-shelf components. Finally it dramatically reduces space requirements and is easily portable. The dimensions of the da Vinci S/Si console are 166 cm x 163 cm x 163 cm and it weighs 273 kg, compared to our setup including 2 HTCVive tracker (7cm × 8cm × 5cm, 100g), 2 basestations (9cm x 6cm x 8cm 0.45 kg ), 2 SCOPEYE wireless input transmitter (in total 39cm x 30cm x 16cm) two sensory gloves (worn by the surgeon, 138g) and a computer. It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field. Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
Future work includes implementing velocity control to enhance the user experience, enabling full control of both surgical arms, upgrading hardware to improve tracking precision, and expanding the repertoire of user gestures to increase the range of available actions. We will also implement enhanced ergonomic features, such as adjustable armrests and customizable control environments.